How do you solve these problems?
x+1/y=7
7n-8m=4-2m

Answer by stanbon(75887)

solve these problems?
x+1/y=7
x + (1/y) = 7
1/y = 7-x
----
Invert::
y = 1/(7-x)
-------------------
7n-8m=4-2m
------
6m = 7n-4
----
m = (7n-4)/6
---------------
Cheers,
Stan H.
----------------